Problem 5-3 Calculating Present Values [LO2] For each of the following, compute the present value (Do not round intermediate calculations and round your final answers to 2 decimal places. (e.g., 32.16): 13 4 29 40 7% 13 14 $16,051 57,557 892,073 556,164
Answer #1
The requirement is to find PV given FV.
PV = FV/(1+r)^n, where r = interest rate and
n = number of years.
1) =16051/1.07^ 13 = $     6,660.59
2) = 57557/1.13^4 = $   35,300.79
3) = 892073/1.14^29 = $   19,959.96
4) = 556164/1.09^40 = $   17,706.92